OpenStudy (anonymous):

Find f if f''(x)=2e^t+3sin(t), f(0)=7, f(pi)=-8

OpenStudy (firekat97):

because you have been given f"(x) you need to integrate twice to get to f'(x) and then f(x)

OpenStudy (anonymous):

I have been doing that but have not gotten the correct response

OpenStudy (firekat97):

Can you post your working so far?

OpenStudy (anonymous):

I have quite a few different problems But I only have one more chance to get it right.

OpenStudy (anonymous):

Like I have worked them a lot

OpenStudy (firekat97):

but can you show me how you went about trying to solve the problem? So your working out so I can see where you're going wrong..

OpenStudy (anonymous):

I am not exactly sure how to take a photo and attach a file since I am on my computer.

OpenStudy (firekat97):

do you think you could type up your solution?

OpenStudy (anonymous):

its a kinda complicated solution thats why i am on here trying to get an answer for it so i an compare the correct answer to my answer to see where i went wrong

OpenStudy (firekat97):

okay so when you integrate 2e^x what do you get?

OpenStudy (anonymous):

2e^x anything e^x is still e^x

OpenStudy (firekat97):

yup thats correct

OpenStudy (firekat97):

sorry what do you get when you integrate 3sint

OpenStudy (anonymous):

-3cost

OpenStudy (firekat97):

Yup, thats correct too, so we know that the integral of f"(x) is 2e^t - 3cost + c

OpenStudy (firekat97):

so f'(x) = 2e^t - 3cost so now we have to integrate again

OpenStudy (anonymous):

I am running out of time its an online homework due at a specific time Can you tell me the answer? I only have one last attempt and 2/23 problems left that I have been struggling with for the past hour and a half

OpenStudy (firekat97):

you're not going to learn if i just tell you the answer..

OpenStudy (firekat97):

just try integrating f'(x) = 2e^t - 3cost + c in the same way and tell me what you get

OpenStudy (anonymous):

The tricky part with this question is that you have two constants since you have to integrate twice I understand the mechanics of how to do the problem and with you not telling me your answer I can't know if it is correct so if the answer you have is wrong then I am learning the wrong thing I have tried 4 different answers so if you tell me your answer I can check to see if that was one of the answers I got

OpenStudy (anonymous):

2e^t+3sint+cx+c

OpenStudy (freckles):

he and you are both correct and saying f'(x)=2e^t-3cos(t)+c you can integrate first... or you can use your constraint to find c...wait isn't one of your constraints written incorrectly...one of them should be f'(a)=b where a and b are some numbers ...whichever one it is use the constraint then integrate note: if you integrate first make sure you don't use the same letter for the other constant you get after the next integration

OpenStudy (anonymous):

But what is the final answer that you get ?

OpenStudy (freckles):

No one can know the problem is incorrectly stated

OpenStudy (anonymous):

What do you mean it is incorrectly stated ?

OpenStudy (freckles):

As I was saying earlier isn't one of your constraints is suppose to be in the form f'(a)=b where and b are some numbers?

OpenStudy (anonymous):

yes it is supposed to be but the two questions I was stuck on do not give it in the form of f(x) and f'(x) which is why it is difficult But it can be done I know the steps but keep getting the wrong answer

OpenStudy (freckles):

where a and b are some numbers*

OpenStudy (firekat97):

No, not necessarily.

OpenStudy (freckles):

err... do you not have a condition on f'?

OpenStudy (firekat97):

In this case, because f(0) is given, cx can be eliminated giving the constant d

OpenStudy (freckles):

you have f(0)=7 and f(pi)=-8 one of these is really suppose to be about f'

OpenStudy (firekat97):

what I mean is f(x) = 2e^t - 3sint + cx + d we can find d and hence when we sub in f(pi) = -8, the value of d is already known so we can find c as well

OpenStudy (anonymous):

no you are supposed to integrate twice fine the constant of the first integrated equation integrate again and then solve for the second constant and plug everything back into the integrated problem that gave the equation for f(x)

OpenStudy (firekat97):

because since f(0) = 7 is known, everything except d is equal to 0, so d = 7

OpenStudy (anonymous):

FireKat97 that is what I have been doing and keep getting the wrong answer Can you tell me your answer so that I can see if it is correct and then send me a picture of your work ?

OpenStudy (firekat97):

then we know that f(pi) = -8 = 2e^π - 3sinπ + cπ + 7 and c can now be found too

OpenStudy (firekat97):

I got f(x) = 2e^t - 3sin(t) + t(-15 - 2e^t)/π + 7 hoping I didn't make any silly mistakes along the way...

OpenStudy (freckles):

I guess you guys were right about f(0)=7 and f(pi)=-8 I was just thinking they normally give a condition for each integration step but I got a different d and c above using the conditions

OpenStudy (firekat97):

Unless f(0) is given, something about f'(x) is required

OpenStudy (freckles):

right I see that now

OpenStudy (anonymous):

the answer you gave was incorrect FireKat97

OpenStudy (firekat97):

do they need it in decimals if its an online quiz thing

OpenStudy (freckles):

but yeah f(0)=7 doesn't give d=7 you should get 2-3(0)+c(0)+d=7 which this will give us a different c in the end also

OpenStudy (firekat97):

OH wait whoops! yeah

OpenStudy (anonymous):

no, no decimals

OpenStudy (anonymous):

so wait what should the answer be ?

OpenStudy (firekat97):

because e^0 = 1 not 0

OpenStudy (firekat97):

d = 5 my badddd

OpenStudy (anonymous):

so instead of 7 it should be 5 ?

OpenStudy (firekat97):

yeah and so when you sub d in again for f(pi) = -8, use d as 5, so c should be different too

OpenStudy (anonymous):

okay i give up I just won't get points for this problem lol

OpenStudy (firekat97):

don't give up, you're almost there :p

zepdrix (zepdrix):

broski, ask for help on a question sooner than 30 minutes before the due date! XD that's crazy sauce!

OpenStudy (anonymous):

I still don't have an answer lol so I do not feel almost there

OpenStudy (anonymous):

I have more than 30 minutes..

zepdrix (zepdrix):

For future reference, when you get into a jam like this, https://www.wolframalpha.com/input/?i=f%27%27%28t%29%3D2e%5Et%2B3sin%28t%29%2C+f%280%29%3D7%2C+f%28pi%29%3D-8 Wolfram is really helpful with problems such as this one. Unless your teacher considers that cheating of course :p I find it to be a useful tool to check work though.

zepdrix (zepdrix):

yay \c:/

OpenStudy (anonymous):

I have already tried wolfram and it does not work either not for this problem

OpenStudy (firekat97):

i copied and pasted and forgot to get fix the 5 :p haha

zepdrix (zepdrix):

inputting into wolfram is a little tricky :O I think I did it correctly though, didn't i?

OpenStudy (anonymous):

so that is not the correct answer FIREKAT?

OpenStudy (freckles):

@Asiah321 we said d=5 and you had f(t)= 2e^t - 3sint + cx + 5 so now use this other condition f(pi) = -8 plug in pi for t and replace f(pi) wth -8 and solve for c

OpenStudy (freckles):

oops that x should be t but whatever

OpenStudy (freckles):

solve for c: -8=2e^pi-3sin(pi)+c(pi)+5

OpenStudy (anonymous):

Where is the typo ?

OpenStudy (jhannybean):

\[f''(x)=2e^t+3\sin(t)\]\[f'(x)=\int (2e^t+3\sin(t))dt = 2e^t-3\cos(t)+c\]\[f(x) = \int (2e^t-3\cos(t)+C)dt = 2e^t-3\sin(t) +Cx+D\]\[7=2e^{0}-3\sin(0)+C(0)+D \implies D = 5\]\[-8 = 2e^{\pi} -3\sin(\pi)+\pi C+5\]\[\pi C = -2e^{\pi}-13 \qquad \implies C=\frac{-2e^{\pi}-13}{\pi}\]\[f(x) = ....\]

OpenStudy (jhannybean):

Therefore, what im getting for \(f(x)\) is.. \[f(x) = 2e^{t}-3\sin(t)+\left(\frac{-2e^{\pi}-13}{\pi}\right)t+5\]

OpenStudy (jhannybean):

some variable errors. But you can figure out that much

OpenStudy (freckles):

lol is it f(x) or f(t)

OpenStudy (jhannybean):

Haha, it's f(t).

OpenStudy (anonymous):

Jhannybean it was correct !!!! thank you so muchhhhh !!!!

OpenStudy (jhannybean):

Whatever, it's f(o)

OpenStudy (freckles):

I only wanted to point that out because there has been a bunch of mixing of variables in this post

OpenStudy (anonymous):

jhannybean can you check your messages please

OpenStudy (jhannybean):

Yeah, that's why i just went with what I thought it was.

OpenStudy (jhannybean):

You shouldn't be thanking me, honestly. I was working it out while @freckles and @FireKat97 we're helping you solve it:P Sometimes you have to work with people instead of asking fr the answer.

OpenStudy (anonymous):

I really appreciate you all ! I have been struggling with this problem and another one for a while

OpenStudy (jhannybean):

post your new question in a new thread :)